# PSAppendix - Selected Primary Studies

### v1

@@ 7. Future Research Directions
Based on the MLOps practices, adoption challenges and proposed solutions, we identify the future research directions to facilitate MLOps adoption. To provide a concise view of our findings, Fig. 4 presents a mapping of MLOps best practices, adoption challenges and their related solutions from RQ2 and RQ3, along with future research directions derived based on the existing gaps.
Figure 4. Mapping of MLOps best practices, adoption challenges, solutions and Future research directions to improve MLOps adoption
@@
Open source tool integration support. To meet different business requirements, customization of MLOps platforms is required where practitioners attempt to integrate open-source tools with managed MLOps platforms [108]. In this case, interoperability is achieved through standardization of interfaces, model and data formats, protocols, etc. Moreover, integration of open-source tools can cause security risks, requiring analysis before integration and continuous audits to detect and mitigate any risks after integration.
@@
MLOps tools for responsible AI. As highlighted in both academic and gray literature such as [152], [105], [164], responsible AI tools (i.e., bias detection and mitigation, provenance extraction, compliance checking) need improvement. Future work on the responsible AI tools requires input from domain experts from different domains having various levels of security, fairness, regulatory, governance requirements that span across the ML life cycle.
@@
Conjugated artifact management solutions. With the increased types and number of artifacts generated during MLOps productionization, conjugated artifact management solutions are required to track the artifacts while maintaining relationships among them. Future solutions should also focus on governance along with efficient and scalable storage of artifacts across distributed environments.
@@ 8. Conclusion
The complexity of ML solutions due to the experimental nature of data and ML, the dynamic nature of production environments, the availability of continuous streams of new data and decentralized computing infrastructure (i.e., edge-cloud integrated environments) makes productionalization of ML models extremely challenging for the practitioners. Hence, MLOps introduces practices and technologies to create continuous integration, continuous delivery, continuous training and continuous monitoring-related autonomous workflows for the smooth transition of ML models from development to deployment, with increased transparency, model quality, and responsible and secure AI support. This has made MLOps a popular topic among practitioners and academia. However, with the innate complexity of MLOps further enhanced by rapid contributions made to MLOps from industry and academia, it is challenging to navigate the current landscape of MLOps.
We have systematically selected and analyzed 150 relevant academic studies and 48 gray literature to provide a comprehensive overview of the current MLOps landscape by addressing three research questions: 1) How do researchers and practitioners perceive MLOps, 2) What are the enabling solutions and best practices proposed for MlOps?, 3) What are challenges practitioners face when adopting MLOps?. We derived future research directions based on the comprehensive analysis of MLOps adoption challenges and existing solutions to drive future research in MLOps. Practitioners and researchers can utilize rigorous analysis and insights of MLOps provided in this work to guide the adoption of MLOps into their ML-enabled application productionalization process with increased understanding of cultural, technical and socio-technical aspects of MLOps and drive future advances to improve MLOps adoption.

### v2

@@ 5.3.5. Migration to MLOps Platforms
Practitioners face significant challenges across socio-technical shifts and barriers, ML pipeline complexity and fragmentation, scalability issues, distributed computing environments, the need for context-specific adaptations, ensuring compliance, fairness, and security of ML systems, and decision dilemmas about migrating to MLOps platforms and balancing automation with human involvement. Proposed solutions to these challenges reflect a shift towards human-centric, systematic, and transparent communication strategies, emphasizing domain-specific architectures, enhanced governance, systematic artifact management, and structured human-in-the-loop interventions.
@@ 6. Future Research Directions
Based on the MLOps practices, adoption challenges and proposed solutions, we identify the future research directions to facilitate MLOps adoption.
@@
Open source tool integration support. To meet different business requirements, customization of MLOps platforms is required where practitioners attempt to integrate open-source tools with managed MLOps platforms [108]. In this case, interoperability is achieved through standardization of interfaces, model and data formats, protocols, etc. Moreover, integration of open-source tools can cause security risks, requiring analysis before integration and continuous audits to detect and mitigate any risks after integration [108], [105].
@@
MLOps tools for responsible AI. As highlighted in both academic and gray literature such as [152], [105], [164], responsible AI tools (i.e., bias detection and mitigation, provenance extraction, compliance checking) need improvement. Future work on responsible AI tools requires input from experts across different domains having various levels of security, fairness, regulatory, governance requirements that span across the ML life cycle.
@@
Conjugated artifact management solutions. With the increased types and number of artifacts generated during MLOps productionization, conjugated artifact management solutions are required to track the artifacts while maintaining relationships among them [122]. Future solutions should also focus on governance along with efficient and scalable storage of artifacts across distributed environments.
7. Related Literature Reviews
We examine existing systematic review studies related to MLOps and qualitatively compare them with our work based on the type of considered primary studies, the research questions covered, the number of primary studies included, and the time period for study selection. We identified eight related studies: three Multivocal Literature Reviews (MLRs), which collected data from both academic and gray literature, and five Systematic Literature Reviews (SLRs), which relied solely on academic literature. Table 3 presents a comparison of these reviews with our work, highlighting our contributions.
Table 3. Comparison of Related Work
Study Is multivocal MLOps Conceptualization MLOps Activities MLOps Roles & Responsibilities Practices & Tools Challenges No. Primary Studies Timeframe
[…]
(Lima et al., 2022) No $Circle \Circle \Circle$ 30 A Up to Jul, 2021
[…]
Our work Yes $CIRCLE \CIRCLE \CIRCLE \CIRCLE$ 150 A, 48 G 2013 - Sep, 2023
[…]
A: Academic literature study, G: gray literature study
Kreuzberger et al. (Kreuzberger et al., 2023) reported a mixed-method research, where they conducted an interview with practitioners in addition to an SLR. Their work targeted one RQ: “what is MLOps”, which they answered with an overview of MLOps activities, roles and responsibilities and used these insights to create a unifying definition for MLOps. While they also discussed several open challenges related to adopting MLOps, it is not the main objective of the study.
Lima et al. (Lima et al., 2022) reported an SLR, where their RQs aimed to identify activities related to ML model deployment, MLOps related roles and responsibilities, operationalization tools for ML models, and challenges in ML model deployment. However, this study lacks comprehensive data synthesis to provide indepth analysis of roles and responsibilities. Additionally, their discussion on MLOps related technologies is limited to MLOps platforms extracted from academic literature and their discussion on challenges does not elaborate on the challenges related to MLOps adoption.
Kolltveit et al. (Kolltveit and Li, 2022) reported an SLR on the operationalization of ML systems, which covers all steps following model training and evaluation. Specifically, they answered four RQs focused on tooling and infrastructure related aspects of MLOps such as; how are ML models operationalized, what are the main challenges in operationalizing ML models, what tools and software infrastructure are used to operationalize ML models, what are the feature gaps in the tooling and infrastructure used to operationalize ML models. However, since the focus is only on tooling and infrastructure for operationalizing ML models, this work does not provide a comprehensive view of MLOps best practices considering the entire ML lifecycle.
Faubel et al. (Faubel and Schmid, 2023) reported an SLR on the technical overview of MLOps. Specifically, the study presented MLOps definitions and listed the most common activities, and listed tools for automating activities in MLOps. However, this review does not consider roles and task distributions in MLOps, thus lacking emphasis on socio-technical aspects in MLOps.
Mboweni et al. (Mboweni et al., 2022) reported an SLR on the state-of-the-art DevOps practices in ML-enabled systems. Specifically, this work aimed to report how MLOps is understood and differs from DevOps. The reported content provides a higher-level overview of MLOps definitions, presents key roles of MLOps and associated tasks, and describes MLOps activities by constructing an MLOps architecture covering ML lifecycle.
Steidl et al. (Steidl et al., 2023) reported a mixed-method investigation (i.e., SLR and a semi-structured review study) focusing specifically on continuous development pipelines of AI models, which is one aspect of MLOps practices. This work focused on analyzing definitions of various terms related to continuous development pipelines for AI such as DevOps for AI, CI/CD for AI, MLOPs, etc., tasks handled by these pipelines, and the challenges related to implementation, adaption and usage of pipelines for continuous development of AI.
John et al. (John et al., 2021) delivered an MLR on the adoption of MLOps in practice. More specifically, the authors derived a framework for MLOps that shows the activities that take part in the continuous development of ML models and maps the practices of companies to MLOps maturity model to show how practices evolve through stages. However, they provided a high-level picture rather than elaborating on details regarding these questions.
Lwakatare et al. (Lwakatare et al., 2020) presented an MLR on the RQ of “how to enable modern software development process with continuous delivery for AI-enabled systems”. The authors considered one academic and eight gray literature items. The study discusses tools and techniques used in the continuous delivery of ML systems. Important details are available concerning the development stages of ML-enabled systems.
Our analysis of the secondary studies presented in Table 3 demonstrates that no existing work contributes to building a holistic and timely depiction of MLOps landscape by answering all three research questions. Also, the majority of the related work relies solely on academic literature (Kreuzberger et al., 2023; Lima et al., 2022; Kolltveit and Li, 2022; Faubel and Schmid, 2023; Mboweni et al., 2022). Given that MLOps is a rapidly evolving, practice-oriented field, insights from gray literature sources such as industry white papers and company websites offer valuable and timely perspectives on how MLOps is perceived by the community and adopted in real-life settings. Furthermore, the rapid improvements in MLOps call for a systematic review of the most recent literature to provide a timely view of the MLOps landscape. However, coverage by existing reviews extends only until 2022.
Considering these, our MLR aims to provide a comprehensive conceptualization of MLOps by integrating insights from a broader range of primary sources, from 150 academic studies with more practical knowledge from 48 gray literature items published up to September 2023. Our study extends prior research by linking MLOps conceptualization to productionalization challenges, identifying emerging activities, roles, practices, challenges, and solutions. Furthermore, we discuss future directions, emphasizing the need for domain-specific frameworks, secure pipelines, automation impact analysis, and responsible AI tools, aiming to enhance MLOps efficiency.
7.1. Evolution of MLOps
We share common RQs with related work. Hence, we compare our findings with those reported in prior studies to illustrate how MLOps has progressed over time.
Conceptualization: The conceptualization of MLOps in earlier systematic studies aligns closely with our RQ1.1 results (e.g., (Kreuzberger et al., 2023), (Mboweni et al., 2022)). This indicates that MLOps definitions in the literature have remained largely consistent from 2022 (the latest timeframe in prior studies) through September 2023 (our study’s timeframe). Also, earlier studies explicitly defined MLOps to clarify its conceptual boundaries (e.g., [78]), whereas recent papers directly use the term without definitions, assuming community familiarity (e.g., [188], [144], [145]). This shift highlights the growing maturity and widespread acceptance of MLOps as a recognized research area.
Activities: While our findings align with previous works regarding core MLOps activities, we identify newer focal areas such as security [105], expanded quality assurance which includes data and model aspects [191], compliance and governance processes [106], and enhanced collaboration and communication management [151]. The increased prominence of these aspects in recent literature signals a maturation in the understanding and structuring of MLOps practices (Section 3.2).
Roles: Recent literature emphasizes new roles focusing on quality assurance, compliance, and security concerns, such as data stewards, data quality boards, and compliance teams [106], [114]. Although an earlier study identified the MLOps engineer role (Kreuzberger et al., 2023), our findings explicitly differentiate its responsibilities from those of ML engineers, positioning it within clearly defined workflows and clarifying its interactions with other emerging roles (Section 3.3).
Practices and techniques for implementing MLOps: Prior systematic studies identified practices like workflow automation, artifact tracking, collaboration, and continuous monitoring. Our findings extend these set of practices by highlighting recent developments, including the use of domain-specific reference architectures for specific business contexts [134], pipeline optimization strategies to enhance scalability and performance [136], containerization and service-oriented architectures [132], and systematization of regulated ML tasks to ensure compliance with emerging governance requirements [117] (Section 4).
MLOps adoption challenges and solutions: Challenges such as ML pipeline complexity, socio-technical barriers, scalability and performance issues, distributed computing environments, and adapting MLOps to specific contexts have been reported in related work. Beyond these, our study highlights more recent and nuanced issues such as compliance [111], fairness [152], and security of ML systems [105], customization of ML platforms [108], pipeline fragmentation due to tool diversity (section 5.1.2), and decision-making challenges regarding automation and migration (sections 5.2.5, 5.3.5). We observe that solutions to these challenges have correspondingly evolved, increasingly incorporating human-centric, transparent, and plan-focused approaches for cultural and cross-silo adaptations [103], domain-specific MLOps architectures, frameworks supporting compliance, fairness, and security, as well as human-in-the-loop interventions [110]. (Section 5).
@@ 8. Conclusion
We have systematically selected and analyzed 150 relevant academic and 48 gray literature studies to provide a comprehensive overview of the current MLOps landscape by addressing three main research questions: 1) Conceptualization of MLOps, 2) What are the state-of-the-art best practices and techniques for implementing MLOps?, 3) What are the challenges practitioners face when adopting MLOps, and what solutions have been proposed to address them? Our findings emphasize emerging aspects of MLOps, i.e., compliance, governance, security, and quality, challenges and practices such as pipeline complexity, domain specific needs, automation trade-offs, and scalability in distributed environments. With these, MLOps practices, roles, and activities evolved toward identifying new roles, collaboration and communication oriented practices, and frameworks that address new emerging and complex aspects of MLOps such as customization and complex infrastructures. We hope this study serves as a valuable resource for researchers and practitioners, guiding future advancements in MLOps adoption and innovation.